Shift Summary for Owl Nov. 7, 1999

S. Elizabeth Turner (eturner@cooper.physics.ucla.edu)
Sun, 7 Nov 1999 05:18:50 -0800 (PST)

Shift Summary for DAY shift, 11/06/99 (Sat.)

On Shift: Albert, Elizabeth

Hours of beam available at all intensities: 3h 05m
Hours of beam available at nominal intensity: 2h 45m
Hours of E799 data (special runs not included): 2h 45 m
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

Blown fuse in E3.
========================================================================
Runs ended this shift:

none (14919 in progress)
========================================================================
Accesses during shift:

None.
========================================================================
Overview of shift:

First half of shift spent waiting for beam. Beam down due to a
blown fuse in E3. After 4:00 when TeV begins running on 85 s
supercycle, very smooth 799 data taking.
========================================================================
Description of Shift:

Run Spill Time Note
--- ----- ---- ----
14919 00:00 Take over from eve shift. No beam. Run
is in progress, but no data has been taken
so far.
00:30 SET adjusts trigger source timing for RCs.
308 03:45 Low intensity beam returns.
314 Intensity above 2E12.
334 04:25 Intensity up to 8.7E12.
371-374 Left Bends trip.
416-418 Left Bends trip.
420-422 Left Bends trip.
========================================================================
Follow up Notes on Previous Shift

Still no B16 triggers.

========================================================================
Items to Follow up on Next Shifts